Circular No. 8324

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V2574 OPHIUCHI = POSSIBLE NOVA IN OPHIUCHUS N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V2574 Oph has been given to this nova (cf. IAUC 8323). Additional magnitudes of V2574 Oph: Apr. 11.38 UT, [13.5: (ASAS-3 image; via H. Yamaoka, Kyushu University); 14.780 and 14.785, 10.9 (K. Haseda, Aichi, Japan; photographic; prediscovery; via Yamaoka); 14.782, 10.5 (Y. Nakamura, Mie, Japan; photographic; independent discovery forwarded by J. Watanabe, National Astronomical Observatory of Japan); 15.785, ∼ 10.5 (A. Takao, Kitakyushu, Japan; via Yamaoka); 16.146, 10.4 (P. Schmeer, Bischmisheim, Germany; visual); 16.403, V = 10.29 (J. D. West, Mulvane, KS); 16.409, B = 11.74 (West); 16.417, R_{c} = 9.52 (West); 16.803, 10.7 (A. Pearce, Nedlands, W. Australia; visual). Yamaoka adds that his measurement of the position of V2574 Oph from an R-band CCD image taken by B. Monard on Apr. 16.144 yields α = 17^{h}38^{m}45^{s}.51, δ = −23^{o}28′18′′.6 (equinox 2000.0; 95 UCAC2 reference stars).

H. E. Bond, Space Telescope Science Institute; and F. Walter, State University of New York at Stony Brook, report that spectrograms (range 350–530 nm; resolution 0.4 nm) of V2574 Oph were obtained by A. Pasten with the SMARTS Consortium 1.5-m telescope at Cerro Tololo Interamerican Observatory on Apr. 16.25–16.27 UT. The spectrum resembles that of an F supergiant, with strong absorption lines of Fe II and no obvious emission lines. The absorption lines are blueshifted by ∼ −410 km/s. The object thus appears to be a slow classical nova caught near maximum light.

SUPERNOVAE 2004bf AND 2004bi T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center

for Astrophysics, write that a spectrogram of SN 2004bf (cf. IAUC 8317), obtained by J. Huchra with the Mt. Hopkins 1.5-m telescope on Apr. 13.42 UT, shows it to be a type-Ic supernova, similar to SN 1987M (see Filippenko et al. 1990, A.J. 100, 1575). The spectrum is reddened and shows a strong, narrow Na I D absorption at the velocity of the host galaxy with an equivalent width of 0.2 nm. The dust maps of Schlegel et al. indicate a low value for Galactic reddening of E(B − V) = 0.012. A spectrum of SN 2004bi (cf. IAUC 8321), obtained by Huchra on Apr. 15.22, shows it to be a type-Ic supernova. The spectrum is similar to early spectra of SN 1994I (Filippenko et al. 1995, Ap.J. 450, L11), but there are some peculiarities. An absorption near 625 nm (rest wavelength) may be Si II (rest 635.5 nm), but it is unusually strong compared to typical type-Ic supernovae.

2004 April 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green